OpenStudy (anonymous):

Choose the point on the terminal side of -210°.

OpenStudy (anonymous):

Add in 360° I think.

OpenStudy (anonymous):

Do you know what terminal side is?

OpenStudy (anonymous):

My choices: A. -1/1 B. -1/sqrt3 C. sqrt3/-1 D. -sqrt3/1

OpenStudy (yttrium):

@MandyNeedsHelp , do you know how to graph this function? That will help you a lot.

OpenStudy (anonymous):

I dont, im looking at my notes right now but its all confusing.

OpenStudy (yttrium):

But do you know what a terminal side is?

OpenStudy (anonymous):

The side that the measurement of an angle ends at

OpenStudy (anonymous):

In all of my examples it gives me the terminal side, thats why im confused.

OpenStudy (yttrium):

You're correct about the terminal side. And you can use tangent to know the point, since the formula is tanx = y/x So using tan(-210), you get what?

OpenStudy (anonymous):

tan is opposite/adjacent but im not sure where i find that?

OpenStudy (yttrium):

Yes that is also true, but in coordinate plane you can use the tan=y/x

OpenStudy (anonymous):

I understand, but what is tan in this problem if im supposed to multiply it by -210

OpenStudy (yttrium):

no, you get the tan of -210 to arrive at an answer.

OpenStudy (anonymous):

Im sorry, im confused. Can you just give me the answer and how you got to it? I have other problems i can do on my own.

OpenStudy (yttrium):

-1/sqrt3

OpenStudy (yttrium):

When you graph it, it will also appear as 150 degrees. And with that you can use the reference angle of 30 degrees because it is 30 degrees near the x-axis. As we all know tan30 =1/sqt3 and since the graph is found in the 2nd quadrant, the x value is negative. That is why it became, -1/sqrt3 

OpenStudy (anonymous):

Thank you so much for explaining it, im starting to get a better understanding of this.